OpenStudy (anonymous):

substitution method 15x-10y=-5 15y=4+5x

OpenStudy (anonymous):

patience is a virtue.... (1) 15x-10y=-5 (2) 15y=4+5x From (1) we have: 15x = -5 +10y x = (-5+10y)/15 (1.1) Substitute into (2): 15y=4+5[ (-5+10y)/15 ] 15y = 4 + (-5+10y)/3 15y = (12-5 + 10y)/3 15y = (7 + 10y)/3 45y = 7 + 10y 35y = 7 y = 1/5 ans. Plug into (1.1): x = [ -5+10(1/5) ]/15 x = ( - 5 + 2 )/15 x = -1/5 ans.

OpenStudy (anonymous):

That is correct!

OpenStudy (anonymous):

thanks

OpenStudy (yuki):

solve for y substitute